find the tenth term in the sequance:
13,17,21,25....
thanks for your time!

Answer by Nate(3500)

13 + 4 ~> 17 + 4 ~> 21 + 4 ~> 25 ....
a(x) = a(1) + (x - 1)d
a(10) = 13 + (9)4
a(10) = 13 + 36 = 49